OpenStudy (anonymous):

If anybody understands this, please help me understand because I'm stuck! GI bisects angle DGH so that m angle DGI is x-3 and m angle IGH is 2x-13. Find the value of x.

OpenStudy (anonymous):



OpenStudy (anonymous):

I don't get it at all .-. I'm sorry I'm terrible at any kind of math

OpenStudy (anonymous):

because GI bisects the angle DGH Therefore <DGI=<IGH

OpenStudy (anonymous):

but how do you find the value of x?

OpenStudy (anonymous):

2x-13=x-3 2 x-x=-3+13 x=10

jhonyy9 (jhonyy9):

nice

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

yw